# Gravitational waves from strong first order phase transitions

### v1

@@ IV Results
Figure 2: Selected volume-averaged quantities for the detonation ($\alpha_{\text{n}}=0.67$, $v_{\text{w}}=0.92$, left-hand-side) and deflagration ($\alpha_{\text{n}}=0.50$, $v_{\text{w}}=0.44$, right-hand-side). Top: wall speed estimates $v_{\text{w}}$. The velocity estimate $v_{w,\phi}$ is computed from the ratio of kinetic and gradient energy of the scalar field (purple dotted line), and $v_{w,N}$ through the rate of change of number of sites in broken phase and number of links (purple solid line). The fraction of sites in the broken phase $f_{\text{b}}$ is plotted in olive green, and a dot-dashed vertical green line indicates when 95% of the sites are in the broken phase (this also marks $t_{\text{ref}}$, the reference time for unequal time correlators). Middle: root mean square enthalpy-weighted 4-velocities for compressional $\bar{U}_{\parallel}$ (in red) and vortical $\bar{U}_{\perp}$ (in blue) components, along with their non-relativistic equivalents $\bar{v}_{\parallel,\perp}$. The total RMS weighted 4-velocity $\bar{U}$ is also shown. Bottom: integral scales, for the compressional and vortical modes, defined in Eq. (39). We also display fitted decay and growth indexes ($\zeta,\lambda$, respectively) for compressional kinetic energy and integral scale.
@@ IV.1 Volume-averaged quantities
We begin with the fractional volume in broken phase and the wall velocities in Fig. 2 (top).
@@
In the detonation (left panel) the fraction of broken phase reaches the 95% at a time $t\approx 950/T_{\text{c}}$ with the transition being completed shortly thereafter, at $t\approx 1400/T_{c}$. The wall speed quickly grows to its expected asymptotic limit of $v_{\text{w}}=0.92$, and is within $200/T_{c}$ time units from the start of the simulation. This rise at the end is merely a consequence of velocity estimators breaking down when the transition is nearly completed: for example one of the estimators for $v_{\text{w}}$ divides by the number of links, which by this time will be close to zero.
@@
The deflagration (right-hand-side panel of figure 2) however, is rather different: the velocity starts to rise towards the expected wall speed of $v_{\text{w}}=0.44$, but never reaches it, as reheating of the surrounding plasma slows down bubble wall velocities to $v_{\text{w}}\approx 0.05-0.1$. This slow-down begins at roughly $t\approx 600/T_{c}$ or alternatively, when the fraction of broken sites $f_{B}$ is roughly $5\%$ and stops when the broken fraction is $65-70\%$, as can be seen on Fig. 2. Also shown is the asymptotic wall velocity from one-dimensional spherical droplet simulations of [60]. The two velocity estimates are in excellent agreement.
@@
The transition then proceeds rather slowly (at least when compared to the initial times) and the time when the box is at $95\%$ in the broken phase $t_{\text{ref}}=4460/T_{c}$, with the transition fully completing at the very end of the run $t_{\text{end}}=8000/T_{c}$. This reheating process was first shown in [32] and marks an important difference between deflagrations in weak and strong phase transitions, as it results in a reduction of the efficiency of kinetic energy production and therefore in a suppression of the resulting gravitational wave spectra.
@@
We now turn our attention to fluid velocities. Fig. 2 (middle row) shows projected and total fluid 3-velocities and weighted 4-velocities. Maximum fluid velocities in each of the components of the weighted 4-velocity are given in Table 4, along with the inferred shock and eddy turn-over times, which are defined using $R_{*}$ as the length scale.
@@
We see that in the detonation, the compressional modes dominate, and show a marked decrease from the maximum, decaying by about 40% by the end of the simulation. The vortical velocity is much smaller: at most $U_{\perp}\simeq 0.05$, and decays by about 15% by the end of the simulation. One also sees that the weighted 4-velocity is around 40% larger than the 3-velocity. This is too large a difference to be accounted for by the Lorentz $\gamma$ factor, which is around 1.03 for the maximum RMS 3-velocity. Instead it points to strong correlations between the enthalpy and the velocity. This is a feature of a shock, rather than a sound wave, where they are in antiphase. Along with the strong visual impression of shocks in Fig. 1, we can infer that the detonation flow is shock-dominated,
@@
By contrast, the deflagration generates a significant amount of vorticity. Although when the weighted 4-velocity $U$ peaks, at $t\approx 1500/T_{c}$, the flow is dominantly compressional, vorticity continues to rise with $\bar{U}_{\perp}$ reaching roughly $0.08$ at $t=3000/T_{c}$. Given the initially faster decay rate of $\bar{U}_{\parallel}$ these end up being of similar order of magnitude at the end of the run, with $\bar{U}_{\parallel}\approx 0.11$ and $\bar{U}_{\perp}\approx 0.08$ at $t=8000/T_{c}$. Indeed, the compressional and vortical modes appear to be reaching a constant ratio: after a time around $5000/T_{\text{c}}$, the ratio between $U_{\parallel}/U_{\perp}$ fluctuates in the range 1.4 – 1.5. The ratio is not equipartition, which would be $U_{\perp}\approx\sqrt{2}U_{\parallel}$, but it does suggest some exchange between the components which merits further exploration with longer runs.
@@
The accompanying animation shows that vortical modes are generated by the interactions of compression waves as bubbles collide, particularly as the leading shock crosses the phase boundary into the broken phase. Vorticity can also be generated by shock collision [73]: a possible sign of this effect is the slight growth in the RMS 3-velocity following the completion of the transition in the detonation. The weighted 4-velocity however decreases.
@@
$\alpha_{\text{n}}=0.5$, $v_{w}=0.44$ $0.21$ $0.09$ $4958$ $11865$
Table 4: Maximum values of each component 4-velocities $U_{\parallel,\perp}$ along with the associated shock and eddy formation times, $t_{\text{sh}}$ and $t_{\text{ed}}$, respectively.
@@
where $\Gamma={\bar{w}}/{\bar{\epsilon}_{0}}$, as the enthalpy density is almost constant after the completion of the transition. This can be seen in Fig. 8, where the thermal energy $e_{Q}\equiv 3\bar{w}/4$ is plotted. Note that, as defined, $\Gamma$ is not exactly equal to the adiabatic index of an ultrarelativistic fluid, as total energy $\bar{\epsilon}_{0}$ is divided between kinetic and thermal energy.
@@ IV.2 Power spectra and time decorrelation
In the top panels of Fig. 3 we show power spectra of the enthalpy-weighted 4-velocity components $\mathcal{P}_{U_{\parallel}}$, $\mathcal{P}_{U_{\perp}}$. In the detonation, there is much more power in the compressional modes than in vortical, at all scales, consistent with our previous analysis of fluid average velocities. The deflagration, with much closer average velocities in both components, exhibits a particular range of scales $20\lesssim kR_{*}\lesssim 200$ where there is a similar amount of power in the vortical modes at later times in the range shown. In Fig. 16 of [35] the maximum ratio of vortical to compressional spectral density power reached is $P_{v_{\perp}}/P_{v_{\parallel}}\approx 0.3$ for this deflagration, while we report that the maximum ratio of 3-velocity power is $\mathcal{P}_{v_{\perp}}/\mathcal{P}_{v_{\parallel}}\approx 1.11$ and for enthalpy-weighted 4-velocity power $\mathcal{P}_{U_{\perp}}/\mathcal{P}_{U_{\parallel}}\approx 0.94$. This signals that Higgsless simulations are relatively less efficient at generating power in vortical modes.
@@
The middle panels of 3 show the shear stress power spectra $\mathcal{P}_{\Pi}(k)$. We can easily see that for the detonation the shape of shear stress $\mathcal{P}_{\Pi}(k)$, consistent with the $k^{-3}$ high-$k$ power law expected for a velocity spectrum going as $k^{-1}$ [36, 37]. The deflagration however, has quite a different shape, which can be ascribed to an important contribution from the vortical modes (see Section 5).
The bottom panels show the gravitational wave power spectra $\mathcal{P}_{\text{gw}}(k)$, with a $k^{-3}$ power law in $\mathcal{P}_{\text{gw}}(k)$ in both cases. This suggests that vortical modes are not contributing significantly to the resulting gravitational wave signal. Our modelling in Section 5 supports this suggestion.
@@
We will now look at the time decorrelation of $v_{\parallel}$, $v_{\perp}$ and $\Pi$. The resulting decorrelation functions $D_{\parallel}(k)$, $D_{\perp}(k)$ and $D_{\Pi}(k)$ are presented in Fig. 4, in the top, middle and bottom panels, respectively. We show a range of wavenumbers around the inverse integral scales (see Fig. 2).
@@
$\begin{split}D_{\parallel}(k,t,t_{\text{ref}})=&\cos\bigg{(}M_{\parallel}c_{\text{s}}k(t-t_{\text{ref}})\bigg{)}\\ &\times\exp\bigg{(}-\frac{1}{2}\overline{V}_{\!\parallel}^{2}k^{2}(t-t_{\text{ref}})^{2}\bigg{)}\end{split}$ (43)
where $\overline{V}_{\!\parallel}$ can be thought of as a sweeping velocity driving the decorrelation of the compressional component of velocity, and $M_{\parallel}$ can be thought of as a Mach number. We note that this type of decorrelation function has already been seen in other literature, for example in simulations of isotropic compressional turbulence of [74]. The fitted values can be found in the upper half of Table 7. We note that the Mach number parameter $M_{\parallel}$ is slightly larger than one. The compressional modes are organised into shocks and hence can travel slightly faster than $c_{\text{s}}$.
@@
$\alpha=0.5$, $v_{w}=0.44$ $0.080$ $0.050$
Table 7: Averages of the root mean square 3-velocities at the UETC reference time $t_{\text{ref}}$ and the end of the simulation $t_{\text{end}}$, for compressional and vortical modes ($\parallel$,$\perp$ , ⟂), divided by $\sqrt{3}$ to give the average in each vector component. The upper row gives values for the detonation, the lower row for the deflagration.
@@
$D_{\perp}(q,t_{1},t_{2})$ $=$ $\exp\left(-\frac{1}{2}\overline{V}_{\!\perp}^{2}q^{2}(t_{1}-t_{2})^{2}\right.$ (44)
-- $\left.\frac{1}{2}A_{\perp}^{2}\sin^{2}(N_{\perp}c_{\text{s}}q(t_{1}-t_{2}))\right).$
@@
The decorrelation of the shear stress at wavenumber $k$ in the deflagration is much less oscillatory that that of the compressional velocity, even though the flow is dominated by the compressional modes. We can understand this as a consequence of different wavenumbers in the velocity field contributing to the decorrelation, leading to a destructive interference. The decorrelation time is around $\pi/c_{\text{s}}k$. In the deflagration, the characteristic Gaussian form of vortical mode decorrelation is visible at wavenumbers where the velocity power spectrum is more vortical, in the range $20\lesssim kR_{*}\lesssim 200$. The decorrelation time is longer, and similar to that of the vortical velocity mode. At lower wave numbers, where compressional modes dominate, the decorrelation function looks more like that of the detonation.
@@ V.1 Theoretical framework
[…]
@@
In order to compute the shear stress UETC, we assume that ${\mathbf{U}}$ is a Gaussian random vector field [75, 37] with a characteristic length scale $R$, such that the spectral densities can be written
@@
[…]
@@
$C[U_{A},U_{B}](k,t_{1},t_{2})=\bar{R}^{-3}\int\frac{d^{3}q}{(2\pi)^{3}}\Gamma_{AB}(\mu,\tilde{\mu})P_{U_{A}}(q,t_{1},t_{2})P_{U_{B}}(\tilde{q},t_{1},t_{2}),$ (48)
where $\bar{R}=\sqrt{R(t_{1})R(t_{2})}$.
In the convolutions, $\tilde{q}=|\tilde{{\mathbf{q}}}|$ and $\tilde{{\mathbf{q}}}={\mathbf{q}}-{\mathbf{k}}$, and the functions $\Gamma_{AB}$ are geometrical factors from the projector algebra. They take the form
$\Gamma_{\parallel\parallel}=(1-\mu^{2})(1-\tilde{\mu}^{2}),$ (49)
$\Gamma_{\perp\perp}=(1+\mu^{2})(1+\tilde{\mu}^{2}),$ (50)
$\Gamma_{\parallel\perp}=(1-\mu^{2})(1+\tilde{\mu}^{2}).$ (51)
@@
$R(t)=R_{*}[1+(\Delta t/t_{\text{d}})]^{\lambda},\quad\bar{U}^{2}_{A}(t)=\bar{U}^{2}_{A,*}[1+(\Delta t/t_{\text{d}})]^{-\zeta}$ (54)
@@
$\bar{t}=\sqrt{t^{\prime}_{1}t^{\prime}_{2}},\quad r=t^{\prime}_{2}/t^{\prime}_{1}.$ (55)
@@
Returning to the gravitational wave power spectrum, and changing integration variables to $\bar{t}$ and $r$, we have that the change in the power spectrum between times $t_{\text{i}}$ and $t$, both assumed much greater than $t_{\text{d}}$, is
@@
$C_{\Delta}[U_{A},U_{B}](k,\bar{t})=\bar{R}^{-3}\int\frac{d^{3}q}{(2\pi)^{3}}\;\Gamma_{AB}(\mu,\tilde{\mu})\Delta_{AB}(q,\tilde{q},k)\bar{P}_{U_{A}}(q,\bar{t})\bar{P}_{U_{B}}(\tilde{q},\bar{t})$ (60)
@@
$\Delta_{AB}(q,\tilde{q},k)=\frac{\bar{t}}{2\bar{R}}\int_{r_{\text{i}}^{-1}}^{r_{\text{i}}}\frac{dr}{r}\cos[k(t_{1}-t_{2})]D_{A}(q,t_{1},t_{2})D_{B}(\tilde{q},t_{1},t_{2}).$ (61)
@@
$\omega_{AB}$ $=$ $\sqrt{\omega_{A}^{2}+\tilde{\omega}_{B}^{2}}.$ (67)
and $c_{\text{s}}^{\prime}=M_{\parallel}c_{\text{s}}$, as an approximation to the fit results in the previous section. The leading term is the --- - term in the sum over signs in $\Delta_{\parallel\parallel}$ [36, 37, 39, 38], which produces the characteristic sharply rising ($k^{9}$) peak in the power spectrum for long-lived flows. In our simulations, the characteristic scale is initially set by the mean bubble spacing $R_{*}$.
In order to distinguish the gravitational production by the fluid flow after the transition has completed from that produced during the bubble collisions, it is useful to look at the rate of increase in the gravitational wave power spectrum. In numerical simulations, we approximate this with a finite difference, which from (59)
@@
The time dependence of the right hand side of (68), in the wavenumber range where the velocity spectral densities are pure power laws, is through the RMS velocities $\bar{U}_{A}$ and the characteristic length scale $\bar{R}$, which we have assumed in the analysis grow as power laws. Then the dominant term amongst the three convolutions depends on wavenumber $k$, length scale $R$ and RMS 4-velocity $\bar{U}_{\parallel}$ as
@@
Inserting the power-law time dependences of $\bar{R}$ and $\bar{U}_{\parallel}$ from Eq. (54), we have
$\frac{1}{H_{\text{n}}}\frac{\Delta{\mathcal{P}}_{\text{gw}}}{\Delta t}(k,\bar{t})\sim(H_{\text{n}}R_{*})(kR_{*})^{2n-1}(\Gamma\bar{U}_{\parallel,*}^{2})^{2}\left(\frac{\bar{t}}{t_{\text{d}}}\right)^{2n\lambda-2\zeta}.$ (71)
At wavenumbers in the range between the peak wavenumber and the inverse shock width we expect the velocity power spectrum to behave as $k^{-1}$ (i.e. with power law index $n=-1$), characteristic of shocks [43, 34].
Hence, provided $12n\lambda-2\zeta<-$ 1, we expect the gravitational wave power spectrum to approach an asymptote as
@@
Note that the first investigation into kinetic energy decay and emisted gravitational wave power came from [35], and it was conducted using the Higgsless simulations on a wider range of values of phase transition strength $\alpha_{n}$ and wall velocity $v_{w}$. However, the integral scale evolution was not taken into account.
@@
The analyticity requirement of the Fourier transform in infinite volume gives $\beta=4$ [64]. Our simulations do not have good scale separation between the box size and the peak wavelength, and are mildly relativistic, so we do not necessarily expect to see the $\lambda=2/7$ and $\zeta=10/7$ power laws emerging. We do however expect that the parametric dependence of the gravitational wave power spectrum on $R_{*}$ and $\bar{U}_{\parallel,*}$ is correctly captured by these arguments. Recalling that the velocity decay timescale $t_{*}$ is proportional to the shock timescale $t_{\text{sh}}\equiv R_{*}/\bar{U}_{\parallel,*}$, and so

### v2

@@ IV Results
Figure 2: Selected volume-averaged quantities for the detonation ($\alpha_{\text{n}}=0.67$, $v_{\text{w}}=0.92$, left-hand-side) and deflagration ($\alpha_{\text{n}}=0.50$, $v_{\text{w}}=0.44$, right-hand-side). Top: wall speed estimates $v_{\text{w}}$. The velocity estimate $v_{w,\phi}$ is computed from the ratio of kinetic and gradient energy of the scalar field Eq. (35) (purple dotted line), and $v_{w,N}$ through the rate of change of number of sites in broken phase Eq. (36) (purple solid line). The fraction of sites in the broken phase $f_{\text{b}}$ is plotted in olive green, and a dot-dashed vertical green line indicates when 95% of the sites are in the broken phase (this also marks $t_{\text{ref}}$, the reference time for unequal time correlators). Middle: root mean square enthalpy-weighted 4-velocities for compressional $\bar{U}_{\parallel}$ (in red) and vortical $\bar{U}_{\perp}$ (in blue) components, along with their non-relativistic equivalents $\bar{v}_{\parallel,\perp}$. The total RMS weighted 4-velocity $\bar{U}$ is also shown. Bottom: integral scales, for the compressional and vortical modes, defined in Eq. (39). We also display fitted decay and growth indexes ($\zeta,\lambda$, respectively) for compressional kinetic energy and integral scale.
@@ IV.1 Volume-averaged quantities
We begin with the fractional volume in the broken phase and the wall velocities, shown in Fig. 2 (top).
@@
In the detonation (left panel) the fraction of broken phase reaches the 95% at a time $t\approx 950/T_{\text{c}}$ with the transition completing shortly thereafter, at $t\approx 1400/T_{c}$. The wall speed quickly grows to its expected asymptotic limit of $v_{\text{w}}=0.92$ within $200/T_{c}$ time units from the start of the simulation. This rise at the end is merely a consequence of velocity estimators breaking down when the transition is nearly completed: for example the wall velocity estimator Eq. (36) divides by the inferred area of the bubble walls $A$, which by this time will be close to zero.
@@
The deflagration (right-hand-side panel of figure 2) however, is rather different: the velocity starts to rise towards the expected wall speed of $v_{\text{w}}=0.44$, but never reaches it, as reheating of the surrounding plasma slows down bubble wall velocities to $v_{\text{w}}\approx 0.05-0.1$. This slow-down begins at roughly $t\approx 600/T_{c}$ or alternatively, when the fraction of broken sites $f_{B}$ is roughly $5\%$ and stops when the broken fraction is $65-70\%$, as can be seen on Fig. 2. Also shown is the asymptotic wall velocity from one-dimensional spherical droplet simulations of Ref. [38]. The two velocity estimates are in excellent agreement.
@@
The transition then proceeds rather slowly (at least when compared to the initial times) and the time when the box is at $95\%$ in the broken phase $t_{\text{ref}}=4460/T_{c}$, with the transition fully completing at the very end of the run $t_{\text{end}}=8000/T_{c}$. This reheating process was first shown in Ref. [37] and marks an important difference between deflagrations in weak and strong phase transitions, as it results in a reduction of the efficiency of kinetic energy production and therefore in a suppression of the resulting gravitational wave spectra.
@@
We now turn our attention to fluid velocities. Fig. 2 (middle row) shows projected and total fluid 3-velocities and weighted 4-velocities. The maximum fluid velocities for each of the components of the weighted 4-velocity are given in Table 4, along with the inferred shock and eddy turn-over times, which are defined using $R_{*}$ as the length scale.
@@
We see that in the detonation, the compressional modes dominate, and show a marked decrease from the maximum, decaying by about 40% by the end of the simulation. The vortical velocity is much smaller: at most $U_{\perp}\simeq 0.05$, and decays by about 15% by the end of the simulation. One also sees that the weighted 4-velocity is around 40% larger than the 3-velocity. This is too large a difference to be accounted for by the Lorentz $\gamma$ factor, which is around 1.03 for the maximum RMS 3-velocity. Instead it points to strong correlations between the enthalpy and the velocity. This is a feature of a shock, rather than a sound wave, where they are in antiphase. Along with the strong visual impression of shocks in Fig. 1, we can infer that the detonation flow is shock-dominated.
@@
By contrast, the deflagration generates a significant amount of vorticity. The flow is predominantly compressional when the weighted 4-velocity $U$ is at its maximum (at $t\approx 1500/T_{c}$). However, vorticity continues to rise with $\bar{U}_{\perp}$ reaching roughly $0.08$ at $t=3000/T_{c}$. Given the initially faster decay rate of $\bar{U}_{\parallel}$, these end up being of similar magnitude at the end of the run, with $\bar{U}_{\parallel}\approx 0.11$ and $\bar{U}_{\perp}\approx 0.08$ at $t=8000/T_{c}$. Indeed, the compressional and vortical modes appear to be reaching a constant ratio: after a time around $5000/T_{\text{c}}$, the ratio between $U_{\parallel}/U_{\perp}$ fluctuates in the range 1.4 – 1.5. The ratio is not equipartition, which would be $U_{\perp}\approx\sqrt{2}U_{\parallel}$, but it does suggest some exchange between the components which merits further exploration with longer simulations.
@@
The accompanying animation (Ref. [29]) shows that vortical modes are generated by the interactions of compression waves as bubbles collide, particularly as the leading shock crosses the phase boundary into the broken phase. Vorticity can also be generated by shock collisions [75]: a possible sign of this effect is the slight growth in the RMS 3-velocity following the completion of the transition in the detonation. The weighted 4-velocity however decreases.
@@
$\alpha_{\text{n}}=0.50$, $v_{w}=0.44$ $0.21$ $0.09$ $4958$ $11865$
Table 4: Maximum values of each compressional ($U_{\parallel}$) and vortical ($U_{\perp}$) four-velocities, along with the associated shock and eddy formation times ($t_{\text{sh}}$ and $t_{\text{ed}}$, respectively).
@@
where $\Gamma={\bar{w}}/{\bar{\epsilon}_{0}}$, as the enthalpy density is almost constant after the completion of the transition. This is demonstrated in Appendix B, where we plot the thermal energy $e_{Q}\equiv 3\bar{w}/4$. Note that, as defined, $\Gamma$ is not exactly equal to the adiabatic index of an ultrarelativistic fluid, as total energy $\bar{\epsilon}_{0}$ is divided between kinetic and thermal energy.
@@ IV.2 Power spectra and time decorrelation
In the top panels of Fig. 3 we show power spectra of the enthalpy-weighted 4-velocity components $\mathcal{P}_{U_{\parallel}}$, $\mathcal{P}_{U_{\perp}}$. In the detonation, there is much more power in the compressional modes than in the vortical modes. This is the case at all scales, consistent with our previous analysis of fluid average velocities. The deflagration, for which the average velocities in the compressional and vortical components were closer to one another, exhibits a particular range of scales $20\lesssim kR_{*}\lesssim 200$ where there is a comparable amount of power in the compressional and vortical modes at later times in the range shown. In Fig. 16 of Ref. [25] the maximum ratio of vortical to compressional spectral density power reached is $P_{v_{\perp}}/P_{v_{\parallel}}\approx 0.3$ for this deflagration, while we report that the maximum ratio of 3-velocity power is $\mathcal{P}_{v_{\perp}}/\mathcal{P}_{v_{\parallel}}\approx 1.11$ and for enthalpy-weighted 4-velocity power $\mathcal{P}_{U_{\perp}}/\mathcal{P}_{U_{\parallel}}\approx 0.94$. This signals that Higgsless simulations are relatively less efficient at generating power in vortical modes.
@@
The middle panels of Fig. 3 show the shear stress power spectra $\mathcal{P}_{\Pi}(k)$. For the detonation, the shape of shear stress $\mathcal{P}_{\Pi}(k)$ is consistent with the $k^{-3}$ high-$k$ power law expected for a velocity spectrum going as $k^{-1}$ [55, 49]. The deflagration, however, has quite a different shape. This can be ascribed to an important contribution from the vortical modes. The apparent impact of vortical modes on the shear stress (and their absence in the gravitational wave power spectra) will be supported by modelling work in Section V.2.
The bottom panels show the gravitational wave power spectra $\mathcal{P}_{\text{gw}}(k)$, with a $k^{-3}$ power law in $\mathcal{P}_{\text{gw}}(k)$ in both cases. This suggests that vortical modes are not contributing significantly to the resulting gravitational wave signal. Again, our modelling in Section 5 will support this suggestion.
@@
We will now look at the time decorrelation of $_{\parallel}, vv_{\perp}$ and $\Pi$. The resulting decorrelation functions $D_{\parallel}(k)$, $D_{\perp}(k)$ and $D_{\Pi}(k)$ are presented in Fig. 4, in the top, middle and bottom panels, respectively. We show a range of wavenumbers around the inverse integral scales (see Fig. 2).
@@
$D_{\parallel}(k,t,t_{\text{ref}})=\cos\bigg(M_{\parallel}c_{\text{s}}k(t-t_{\text{ref}})\bigg)\\ \times\exp\bigg(-\frac{1}{2}\overline{V}_{\!\parallel}^{2}k^{2}(t-t_{\text{ref}})^{2}\bigg)$ (43)
where $\overline{V}_{\!\parallel}$ can be thought of as a sweeping velocity driving the decorrelation of the compressional component of velocity, and $M_{\parallel}$ can be thought of as a Mach number. We note that this type of decorrelation function has already been seen in other literature, for example in simulations of isotropic compressional turbulence of Ref. [70]. The fitted values can be found in the upper half of Table 7. We note that the Mach number parameter $M_{\parallel}$ is slightly larger than one. The compressional modes are organised into shocks and hence can travel slightly faster than $c_{\text{s}}$.
@@
$\alpha=0.50$, $v_{w}=0.44$ $0.080$ $0.050$
Table 7: Averages of the root mean square 3-velocities at the UETC reference time $t_{\text{ref}}$ and the end of the simulation $t_{\text{end}}$, for compressional and vortical modes ($\parallel,\perp$), divided by $\sqrt{3}$ to give the average in each vector component. The upper row gives values for the detonation, the lower row for the deflagration.
@@
$D_{\perp}(q,t_{1},t_{2})=\exp\left(-\frac{1}{2}\overline{V}_{\!\perp}^{2}q^{2}(t_{1}-t_{2})^{2}\right.\\ -\left.\frac{1}{2}A_{\perp}^{2}\sin^{2}(N_{\perp}c_{\text{s}}q(t_{1}-t_{2}))\right).$ (44)
@@
The decorrelation of the shear stress at wavenumber $k$ in the deflagration is much less oscillatory that that of the compressional velocity, even though the flow is dominated by the compressional modes. We can understand this as a consequence of different wavenumbers in the velocity field contributing to the decorrelation, leading to destructive interference. The decorrelation time is around $\pi/c_{\text{s}}k$. In the deflagration, the characteristic Gaussian form of vortical mode decorrelation is visible at wavenumbers where the velocity power spectrum is more vortical, in the range $20\lesssim kR_{*}\lesssim 200$. The decorrelation time is longer, and similar to that of the vortical velocity mode. At lower wave numbers, where compressional modes dominate, the decorrelation function looks more like that of the detonation.
@@ V.1 Theoretical framework
${\mathcal{P}}_{\text{gw}}(k,t)=\frac{(16\pi G)^{2}}{12H_{\text{n}}^{2}}\\ \times\int_{t_{*}}^{t}dt_{1}\int_{t_{*}}^{t}dt_{2}\dot{G}(k,t,t_{1})\dot{G}(k,t,t_{2})\mathcal{P}_{\Pi}(k,t_{1},t_{2}),$ (45)
@@
In order to compute the shear stress UETC, we assume that $\mathbf{U}$} is a Gaussian random vector field (see Refs. [24, 49]) with a characteristic length scale $R$, such that the spectral densities can be written
@@
where the indices $A,B$ take values $\parallel$ (compressional) and $\perp$ (vortical), and where $\int dx\,x^{2}\tilde{P}_{U_{A}}(x)/2\pi^{2}=1$. The scale can be fixed by requiring that the maximum of $\tilde{P}_{U_{A}}(x)$ is at $x=1$, or some other convenient choice. We will take $R$ to coincide with $R_{*}$ at the time the kinetic energy density reaches its maximum.
@@
$C[U_{A},U_{B}](k,t_{1},t_{2})\\ =\bar{R}^{-3}\int\frac{d^{3}q}{(2\pi)^{3}}\Gamma_{AB}(\mu,\tilde{\mu})P_{U_{A}}(q,t_{1},t_{2})P_{U_{B}}(\tilde{q},t_{1},t_{2}),$ (48)
and $\bar{R}=\sqrt{R(t_{1})R(t_{2})}$.
[…]
$\Gamma_{\parallel\parallel}$ $=(1-\mu^{2})(1-\tilde{\mu}^{2})$ (49)
$\Gamma_{\perp\perp}$ $=(1+\mu^{2})(1+\tilde{\mu}^{2})$ (50)
$\Gamma_{\parallel\perp}$ $=(1-\mu^{2})(1+\tilde{\mu}^{2}),$ (51)
@@
$R(t)$ $=R_{*}[1+(\Delta t/t_{\text{d}})]^{\lambda},$ (54)
$\bar{U}^{2}_{A}(t)$ $=\bar{U}^{2}_{A,*}[1+(\Delta t/t_{\text{d}})]^{-\zeta},$ (55)
@@
$\bar{t}=\sqrt{t^{\prime}_{1}t^{\prime}_{2}},\qquad r=t^{\prime}_{2}/t^{\prime}_{1}.$ (56)
@@
Returning to the gravitational wave power spectrum, and changing integration variables to $\bar{t}$ and $r$, we have that the change in the power spectrum between times $t_{\text{i}}$ and $t$ ($t_{\text{i}}^{\prime}=t_{d}+\Delta t_{i}$ defined analogously to $t^{\prime}$), both assumed to be much greater than $t_{\text{d}}$, is
@@
$C_{\Delta}[U_{A},U_{B}](k,\bar{t})=\bar{R}^{-3}\int\frac{d^{3}q}{(2\pi)^{3}}\;\Gamma_{AB}(\mu,\tilde{\mu})\Delta_{AB}(q,\tilde{q},k)\bar{P}_{U_{A}}(q,\bar{t})\bar{P}_{U_{B}}(\tilde{q},\bar{t}).$ (61)
@@
$\omega_{AB}$ $=\sqrt{\omega_{A}^{2}+\tilde{\omega}_{B}^{2}},$ (68)
and $c_{\text{s}}^{\prime}=M_{\parallel}c_{\text{s}}$, as an approximation to the fit results in the previous section. The leading term is the -- term in the sum over signs in $\Delta_{\parallel\parallel}$ [55, 49, 78, 80], which produces the characteristic sharply rising ($k^{9}$) peak in the power spectrum for long-lived flows. In our simulations, the characteristic scale is initially set by the mean bubble spacing $R_{*}$.
In order to distinguish the gravitational production by the fluid flow after the transition has completed from that produced during the bubble collisions, it is useful to look at the rate of increase in the gravitational wave power spectrum. In numerical simulations, we approximate this with a finite difference, which from Eq. (60) takes the form
@@
The time dependence of the right hand side of Eq. (69), in the wavenumber range where the velocity spectral densities are pure power laws, is through the RMS velocities $\bar{U}_{A}$ and the characteristic length scale $\bar{R}$, which we have assumed in the analysis grow as power laws. Then, the dominant term amongst the three convolutions depends on wavenumber $k$, length scale $R$ and RMS 4-velocity $\bar{U}_{\parallel}$ as
@@
Inserting the power-law time dependences of $\bar{R}$ and $\bar{U}_{\parallel}$ from Eqs. (54, 55), we have
$\frac{1}{H_{\text{n}}}\frac{\Delta{\mathcal{P}}_{\text{gw}}}{\Delta t}(k,\bar{t})\\ \sim(H_{\text{n}}R_{*})(kR_{*})^{2n-1}(\Gamma\bar{U}_{\parallel,*}^{2})^{2}\left(\frac{\bar{t}}{t_{\text{d}}}\right)^{2n\lambda-2\zeta}.$ (72)
At wavenumbers in the range between the peak wavenumber and the inverse shock width we expect the velocity power spectrum to behave as $k^{-1}$ (i.e. with power law index n=1n=-1), characteristic of shocks [39, 40].
Hence, provided $2n\lambda-2\zeta<-1$, we expect the gravitational wave power spectrum to approach an asymptote as
@@
Note that the first investigation into kinetic energy decay and emisted gravitational wave power came from Ref. [25], and it was conducted using the Higgsless simulations on a wider range of values of phase transition strength $\alpha_{n}$ and wall velocity $v_{w}$. However, the integral scale evolution was not taken into account.
@@
The analyticity requirement of the Fourier transform in infinite volume gives $\beta=4$ (see Ref. [43]). Our simulations do not have good scale separation between the box size and the peak wavelength, and are mildly relativistic, so we do not necessarily expect to see the $\lambda=2/7$ and $\zeta=10/7$ power laws emerging. We do however expect that the parametric dependence of the gravitational wave power spectrum on $R_{*}$ and $\bar{U}_{\parallel,*}$ is correctly captured by these arguments. Recalling that the velocity decay timescale $t_{*}$ is proportional to the shock timescale $t_{\text{sh}}\equiv R_{*}/\bar{U}_{\parallel,*}$, and so